# Frustrated Random Walks: A Fast Method to Compute Node Distances on Hypergraphs

### v1

@@ Abstract
A hypergraph is a generalization of a graph that arises naturally when attribute-sharing among entities is considered. Although a hypergraph can be converted into a graph by expanding its hyperedges into fully connected subgraphs, going the reverse way is computationally complex and NP-complete. We therefore hypothesize that a hypergraph contains more information than a graph. In addition, it is more convenient to manipulate a hypergraph directly, rather than expand it into a graph. An open problem in hypergraphs is how to accurately and efficiently calculate their node distances. Estimating node distances enables us to find a node’s nearest neighbors, and perform label propagation on hypergraphs using a K-nearest neighbors (KNN) approach. In this paper, we propose a novel approach based on random walks to achieve label propagation on hypergraphs. We estimate node distances as the expected hitting times of random walks. We note that simple random walks (SRW) cannot accurately describe highly complex real-world hypergraphs, which motivates us to introduce frustrated random walks (FRW) to better describe them. We further benchmark our method against DeepWalk, and show that while the latter can achieve comparable results, FRW has a distinct computational advantage in cases where the number of targets is fairly small. For such cases, we show that FRW runs in significantly shorter time than DeepWalk. Finally, we analyze the time complexity of our method, and show that for large and sparse hypergraphs, the complexity is approximately linear, rendering it superior to the DeepWalk alternative.
@@ I Introduction
A graph is a useful data structure for describing complex relations in real world. In a graph, nodes are connected by edges, and each edge contains exactly two nodes. A hypergraph is an extension of ordinary graphs where each hyperedge of a hypergraph may contain an arbitrary number of nodes (see Fig. 5 in Appendix A for an illustration). In the special case when each hyperedge contains just two nodes, the hypergraph reduces to a graph. Each hyperedge of a hypergraph is considered a clique in a graph. We can therefore convert a hypergraph into a graph by expanding each hyperedge. However, hypergraphs contain more information than what can be encapsulated into its expanded graphs. If we simply expand a hypergraph into a graph, the explicit information of the cliques is lost. Although the conversion from hypergraphs to graphs is reversible in theory, the NP-completeness of detecting maximal cliques in a graphs renders it irreversible in practice. Furthermore, the expansion of hyperedges can be computationally expensive and memory-intensive considering that a hyperedge of $N$ nodes yields $N(N-1)/2$ undirected edges in a graph. It is hence advantageous to directly work with hypergraphs whenever possible rather than with their expanded graphs.
One of the central problems in graph theory is the quantitative determination of graph node distances. The literature has multiple algorithms that aim to achieve this, such as DeepWalkPerozzi et al. (2014), graphSAGEHamilton et al. (2017), hitting times of random walksKlein and Randić (1993) and frustrated random walksLi and Le (2020), among others. We can easily generalize DeepWalk to hypergraphs, yet for the other algorithms, such generalization is highly non-trivial. In this paper, we generalize frustrated random walks to hypergraphs, and show that the approach is on par with the performance of DeepWalk even for complex hypergraphs. There are four reasons that motivate using frustrated random walks (FRW) as opposed to DeepWalk. First, in applications where the goal is to find the nearest neighbors of a few nodes in a large hypergraph, FRW is a preferable and faster option than DeepWalk. Second, FRW conveniently gives a closed-form and interpretable solution for node distances, a solution that is impossible to obtain using any deep-learning based method. Third, FRW does not require any parameter tuning, whereas DeepWalk requires heavy investment in parameter tuning. Finally, and unlike in DeepWalk, the node distances of FRW are asymmetric, rendering it more suitable to describe real-world relationships which are generally asymmetric and non-reflective.
@@
In the above definition, $w$ is the weight of vertex $v_{i}$ in hyperedge $E_{\alpha}$. If the hypergraph is unweighted, then $w=1$ always holds. If we think of $v_{i}$ as a member of a community $E_{\alpha}$, then $e_{i\alpha}$ can be thought of as the loyalty of $v_{i}$ to $E_{\alpha}$. By definition, the degree of a node $v_{i}$ is $D_{i}=\sum_{\alpha}e_{i\alpha}$, and the degree of a hyperedge is $\delta_{\alpha}=\sum_{i}e_{i\alpha}$. We can think of $\delta_{\alpha}$ as the adhesiveness of $E_{\alpha}$.
Researchers have long used random walks to study hypergraphs. In Ref. Zhou et al. (2006), the authors generalized spectral clustering Shi and Malik (2000); Meilă and Shi (2000); Ng et al. (2002) from graphs to hypergraphs and gave their algorithm a random walk interpretation. It is however noted that the hypergraph random walks defined in Ref. Zhou et al. (2006) are no different than the random walks performed on expanded graphs Agarwal et al. (2006); Chitra and Raphael (2019). To take advantage of the higher order structure in hypergraphs, we need to take hyperedge degree into account when performing random walks. Since a hyperedge in a hypergraph represents an adhesive community, it is argued in Carletti et al. (2020) that a random walker roaming on a hypergraph should show preference towards hyperedges of higher degrees (stronger adhesiveness). This is not the only way to generalize random walks from graphs to hypergraphs, but it does demonstrate its advantages, as shown in Ref. Carletti et al. (2020). We also note that in Ref. Carletti et al. (2020), the authors assume that all the hypergraph nodes have the same weight, thus restricting their results only to unweighted hypergraphs. Here in this paper, we will generalize the random walks described in Ref. Carletti et al. (2020) to weighted hypergraphs, and introduce the concept of frustrated random walks on hypergraphs. We show that for heavily-weighted and scale-free hypergraphs, the definition of which will be given later, frustrated random walks are more suitable for computing node distances.
@@
II A unified framework for calculating expected hitting times of random walks
@@
With the abundance of research on random walks on hypergraphs, we note that the vast majority of previous work has focused on the Laplacian matrix. Spectral clustering of eigenvectors of Laplacian matrices Shi and Malik (2000); Ng et al. (2002) naturally leads to image segmentation and community detection in hypergraphs Hayashi et al. (2020); Carletti et al. (2021). Furthermore, the eigenvectors of a Laplacian matrix yield node embeddings which enable us to apply powerful machine learning algorithms to perform node classification on hypergraphs. Since the study of the Laplacian matrix yields information about the long-term stationary distribution of particles that are randomly walking on hypergraphs, we call this approach to studying random walks the stationary approach. On the other hand, we can also extract much information about hypergraphs by focusing on the diffusion process itself, and refer to this approach of studying random walks as being dynamic. In this paper, we will focus on the hitting times of random walks. Given a hypergraph $H$, we select a node $t$ as target, and perform random walks starting from any other node $s$. The hitting time for this process is defined as the number of steps a random walker needs to traverse before it hits the target $t$ for the first time. By definition, the hitting time on a any hypergraph is a random variable that depends on the hypergraph structure, the starting node $s$ and target node $t$, and can thus be denoted as $N^{(s)}_{t}$. The expectation of $N^{(s)}_{t}$ provides a natural measurement of the distance from $s$ to $t$. Note that the distance here is asymmetric, meaning that the distance from $s$ to $t$ is not guaranteed to be identical to that from $t$ to $s$. Thus, the distance here is not the distance function in the mathematical sense, and does not necessarily satisfy all the conditions, especially the symmetricity and triangle inequality, that are required of a mathematical distance function. Since we want to use the expected hitting time to measure the closeness between two persons in the real world, and we know that the real-world human-human relationships are generally asymmetric, non-equivalent, and non-reflexive, this lack of symmetry is thus considered desirable.
@@
The dynamics of random walks on hypergraphs are fully captured by the transition matrix, and different definitions of transition matrices produce different results. We calculate hypergraph node distances to find the nearest neighbors of a node. To accomplish this, we select a node as the target, and calculate the expected hitting times starting from all other nodes. We then rank all those nodes according to their distances to the target. Intuitively, we expect that the shorter a node’s distance to the target, the closer they should be. If we represent real-world human-human interactions using a hypergraph, we expect to find a person’s close friends from among his/her nearest neighbors in the hypergraph. In this paper, we will evaluate the quality of our method’s results by checking whether a person’s close friends show up in his/her top neighbor list.
II.1 Simple random walks on hypergraphs
@@
As already stated, the calculation of hypergraph hitting times requires specification of the random walk’s transition matrix. To capture a random walker’s preference towards a more cohesive hyperedge, we generalize the method in Ref. Carletti et al. (2020) from unweighted to weighted hypergraphs, and define transition probability from node $i$ to node $j$ as
$T^{S}_{ij,i\neq j}=\frac{\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}{\sum_{j}\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}$ (2)
@@
II.2 Frustrated random walks on hypergraphs
Although simple random walk as described in Section II.1 is a good foundational algorithm for exploring hypergraph properties and can be used to generate random paths for DeepWalk, we find that the hitting times derived from this algorithm are not a good measure of node distances when the hypergraph is heavily weighted and scale-freeBarabási and Albert (1999). The fact that most real-world human-human interactions can be recast into a scale-free network compels us to develop an alternative random walk scenario which we call the frustrated random walks algorithmLi and Le (2020), an algorithm that takes into account how people interact with each other in the real world. Real world networks are mostly dominated by the Mathew effect. For example, in a social network, the more followers you have, the easier it is for you to attract new followers, a phenomenon that is called preferential attachmentBarabási and Albert (1999). The advantage of frustrated random walks over simple random walks is that it can yield node distances that are consistent with human judgment even for heavily weighted, scale-free networks. For networks that are lightly weighted or non-scale-free, results from frustrated and simple random walks tend to be similar.

### v2

@@ Abstract
A hypergraph is a generalization of a graph that arises naturally when attribute-sharing among entities is considered. Compared to graphs, hypergraphs have the distinct advantage that they contain explicit communities and are more convenient to manipulate. An open problem in hypergraph research is how to accurately and efficiently calculate node distances on hypergraphs. Estimating node distances enables us to find a node’s nearest neighbors, which has important applications in such areas as recommender system, targeted ads, etc. In this paper, we propose using expected hitting times of random walks to compute hypergraph node distances. We note that simple random walks (SRW) cannot accurately compute node distances on highly complex real-world hypergraphs, which motivates us to introduce frustrated random walks (FRW) for this task. We further benchmark our method against DeepWalk, and show that while the latter can achieve comparable results, FRW has a distinct computational advantage in cases where the number of targets is fairly small. For such cases, we show that FRW runs in significantly shorter time than DeepWalk. Finally, we analyze the time complexity of our method, and show that for large and sparse hypergraphs, the complexity is approximately linear, rendering it superior to the DeepWalk alternative.
@@ I Introduction
A graph is a useful data structure for describing complex relations in real world, such as users on social media, audiences of movies, etcNewman (2018). In a graph, nodes are connected by edges, and each edge connects exactly two nodes. A hypergraph is an extension of ordinary graphs where each hyperedge of a hypergraph may contain an arbitrary number of nodesZhou et al. (2006). In the special case when each hyperedge contains just two nodes, the hypergraph reduces to a graph. Each hyperedge of a hypergraph is considered a clique in a graph. We can therefore convert a hypergraph into a graph by expanding each hyperedge. Fig. 1 illustrates a hypergraph and its corresponding expanded graph.
Figure 1: Panel (a): A hypergraph; panel (b): The expanded graph. The conversion of graph to hypergraph is possible in theory, yet almost impossible in practice.
Hypergraphs arise naturally when there are explicit communities of nodes. In social media, for instance, users are represented as nodes, and users belonging to the same group are collected into the same hyperedge. With the hypergraph representation of users, we want to quantify the similarity between any pair of users, and use this information to suggest friends or recommend products. Intuitively, the similarity between users should be negatively correlated with their distance, thus the computation of hypergraph node distances is a prerequisite for the implementation of recommender systems and targeted adsSarwar et al. (2000); Fouss et al. (2007); Wu et al. (2022).
Although we can easily convert a hypergraph into a graph by expanding its hyperedges, we claim that hypergraphs contain more information than what can be encapsulated into its expanded graphs due to the explicit presence of communities in the former. Furthermore, the expansion of hyperedges can be computationally expensive and memory-intensive. It is thus advantageous to directly work with hypergraphs whenever possible rather than with their expanded graphs.
@@
In the above definition, $w>0$ is the weight of vertex $v_{i}$ in hyperedge $E_{\alpha}$. If the hypergraph is unweighted, then $w=1$ always holds. If we think of $v_{i}$ as a member of a community $E_{\alpha}$, then $e_{i\alpha}$ can be thought of as the loyalty of $v_{i}$ to $E_{\alpha}$. By definition, the degree of a node $v_{i}$ is $D_{i}=\sum_{\alpha}e_{i\alpha}$, and the degree of a hyperedge is $\delta_{\alpha}=\sum_{i}e_{i\alpha}$. We can think of $\delta_{\alpha}$ as the adhesiveness of $E_{\alpha}$.
Given the important applications of computing hypergraph node distances, we have developed a method based on random walks for this task. We will use the expected hitting times of random walks as node distances on hypergraphs. We will show that the simple random walk method is insufficient for the description of complex human-human interactions, and propose our frustrated random walk method. In frustrated random wallks, we augment each transition with an acceptance probability, and thus obtain a different transition matrix from the one in simple random walks. Although the simple and frustrated random walks have different transition matrices, we have developed a unified theoretical framework to efficiently compute their expected hitting times. In summary, the major contributions of our work are as follows:
  1. 1.
    We propose frustrated random walks on hypergraphs, a random walk scenario that outperforms simple random walks for computing node distances in heavily-weighted and scale-free hypergraphs.
  2. 2.
@@
    We design a unified theoretical framework for computing the expected hitting times of random walks, whether it be simple or frustrated, and use the expected hitting times as hypergraph node distances.
  3. 3.
    We test our methods on real-world hypergraph datasets and demonstrate that frustrated random walk method significantly outperforms existing methods in accuracy and/or running speed.
II Background
The importance of the computation of graph node distances has already been widely acknowledgedSommer et al. (2016), with significant applications in social networks (friend suggestion, link prediction, targeted ads), e-commerce (product recommendation), etc. Intuitively, the shorter the distance between two nodes, the more similar they are to each other. Therefore, the primary purpose of computing node distances is to determine node similarities, thus enabling the identification of the nearest neighbors of a target node within a graph. An obvious candidate for computing graph node distances is Dijkstra’s algorithm, which can quickly provide the shortest or geodesic distance from a source node to a target node. Although this algorithm is quick, it focuses solely on the shortest path and ignores all possible alternative routes. However, in a graph representation of social networks, the similarity between two nodes is expected to increase with the number of available alternative routes connecting them. This expectation has led to the development of algorithms that consider the entire graph structure when computing node distances. A plethora of such algorithms, which can be crudely classified into two categories, already exist in the literature. Below, we provide a brief review of these algorithms, focusing only on those most relevant to our work, rather than attempting an exhaustive review of all existing methods.
The first category of algorithms relies on the study of random walks on graphs, and uses expected hitting time (also known as first-passage time) and commute time to evaluate node distancesKlein and Randić (1993); White and Smyth (2003); Fouss et al. (2007); Li and Le (2020). In this framework, for a pair of graph nodes denoted as $v_{i}$ and $v_{j}$, the hitting time $N_{j}^{(i)}$ of a random walk process is defined as the number of steps a random walker starting from node $v_{i}$ needs to traverse before it hits node $v_{j}$ for the first time. $N_{j}^{(i)}$ is a random variable whose expectation value $\mathbb{E}N_{j}^{(i)}$ can be used to measure the distance from $v_{i}$ to $v_{j}$. Here, the distance should be understood as the dissimilarity between nodes $v_{i}$ and $v_{j}$ since it does not satisfy the symmetricity and triangle-inequality conditions that are required of a mathematical distance function. From the expected hitting times, we can construct the expected commute time, defined as $n(i,j)=\mathbb{E}N_{j}^{(i)}+\mathbb{E}N_{i}^{(j)}$, which satisfies all the conditions of a distance function in mathematics, and is therefore a genuine distance in the strict mathematical sense. Given a graph structure, the closed form expressions for $\mathbb{E}N_{j}^{(i)}$ and $n(i,j)$ can be readily obtained from the pseudo-inverse of the graph Laplacian, as is shown in Ref. Fouss et al. (2007). It is however shown in Ref. Li and Le (2020) that for complex and scale-free graphs whose node degree distribution follows a power law, the expected hitting times of simple random walks cannot accurately evaluate node distances, which necessitates the introduction of frustrated random walks to graphs. Here, we will generalize these methods to hypergraphs, and show that frustrated random walks have distinct advantage in measuring node distances of heavily weighted and scale-free hypergraphs.
The second category of algorithms relies on graph node embedding, also called graph representation learning, which means mapping graph nodes to low-dimensional dense vectors and using the cosine distances between these vectors as graph node distancesHamilton et al. (2017a). This includes such algorithms as DeepWalkPerozzi et al. (2014), node2vecGrover and Leskovec (2016), structural deep network embedding (SDNE)Wang et al. (2016), graphSAGEHamilton et al. (2017b), graph attention networksVeličković et al. (2017), etc. These algorithms aim to map graph nodes to a continuous vector space so that nodes that are neighbors in the original graph are mapped to vectors with small cosine distances. There is no universal definition of neighborhood in these algorithms. For example, in DeepWalk, node2vec and graphSAGE, nodes that frequently co-appear in sampled random paths are considered neighbors, whereas in SDNE, neighboring nodes are those that are of first-order or second-order proximity with respect to each other in the original graph. Despite these differences, all algorithms adhere to the guiding principle that neighboring nodes should be mapped to neighboring vectors during the node embedding process. After node embedding, we can use cosine distances between mapped node vectors as graph node distances. Generally, the cosine distances from the node embedding methods is a good metric for evaluating graph node distances, however complex the graphs may be. Among these algorithms, the DeepWalk method excels at measuring node distances of scale-free graphs. A second advantage of DeepWalk in comparison with the other node embedding algorithms is the ease with which it can be generalized to hypergraphs, which prompts us to use it as our benchmark method in our work.
With the abundance of algorithms for computing graph node distances, a natural next step is to generalize them to hypergraphs. The generalization of DeepWalk from graphs to hypergraphs is trivial. To understand why, a brief review of DeepWalk method is necessary. To accomplish graph node embedding using DeepWalk, we first perform random walks on a graph, then record the paths that were generated during the random walk process, interpret the nodes as words and the random walk paths as corpus, and finally invoke word2vecMikolov et al. (2013a, b) to map graph nodes to vectors. Since the generation of random paths in this method is agnostic about whether we are performing random walks on graphs or hypergraphs, a generalization of it to hypergraphs is straightforward.
Although we can easily generalize DeepWalk to hypergraphs, for the other algorithms, such a generalization is highly non-trivial. In this paper, we generalize frustrated random walks to hypergraphs, and show that its performance is on par with the that of DeepWalk even for complex and scale-free hypergraphs. Compared to DeepWalk, frustrated random walks (FRW) has the following four advantages. First, in applications where the goal is to find the nearest neighbors of a few nodes in a large hypergraph, FRW is a preferable and faster option than DeepWalk. Second, FRW conveniently gives a closed-form and interpretable solution for node distances, a solution that is impossible to obtain using traditional deep-learning-based methods. Third, FRW does not require any parameter tuning, whereas DeepWalk requires heavy investment in this task. Fourth, and unlike that in DeepWalk or any other node embedding algorithm, the node distances of FRW are asymmetric, rendering it more suitable to describe real-world relationships which are generally non-equivalent and non-reflective.
Researchers have long used random walks to study hypergraphs. In Ref. Zhou et al. (2006), the authors generalized spectral clustering Shi and Malik (2000); Meilă and Shi (2000); Ng et al. (2002) from graphs to hypergraphs and gave their algorithm a random walk interpretation. It is however noted that the hypergraph random walks defined in Ref. Zhou et al. (2006) are no different from the random walks performed on expanded graphs Agarwal et al. (2006); Chitra and Raphael (2019). To take advantage of the higher order structure in hypergraphs, we need to take hyperedge degree into account when performing random walks. Since a hyperedge in a hypergraph represents an adhesive community, it is argued in Ref. Carletti et al. (2020) that a random walker roaming on a hypergraph should show preference towards hyperedges of higher degrees or stronger adhesiveness. This is not the only way to generalize random walks from graphs to hypergraphs, but it does demonstrate its advantages, as shown in Ref. Carletti et al. (2020). We also note that in Ref. Carletti et al. (2020), the authors assume that all the hypergraph nodes have the same weight, thus restricting their results only to unweighted hypergraphs. Here in this paper, we will generalize the random walks described in Ref. Carletti et al. (2020) to weighted hypergraphs, and introduce the concept of frustrated random walks on hypergraphs. We will show that for heavily-weighted and scale-free hypergraphs, the exact definition of which will be given later, frustrated random walks are more suitable for computing node distances.
Without loss of generality, we will assume throughout the paper that the hypergraphs are connected, meaning that there exists a path between any pair of nodes on a hypergraph. The ease with which we can detect connected components of a hypergraph ensures that this assumption does not compromise the applicability of our method, since we can always apply our method to each connected component of a disconnected hypergraph.
III A unified framework for calculating expected hitting times of random walks
@@
Despite the cornucopia of research on random walks on hypergraphs, the vast majority of previous work has focused on the Laplacian matrix. Spectral clustering of eigenvectors of Laplacian matrices Shi and Malik (2000); Ng et al. (2002) naturally leads to image segmentation and community detection in hypergraphs Hayashi et al. (2020); Carletti et al. (2021). Furthermore, the eigenvectors of a Laplacian matrix yield node embeddings which enable us to apply powerful machine learning algorithms to perform node classification on hypergraphs. Since the study of the Laplacian matrix yields information about the long-term stationary distribution of particles that are randomly roaming on hypergraphs, we call this method of studying random walks the stationary approach. On the other hand, we can also extract precious information about hypergraphs by focusing on the diffusion process itself, and refer to this approach to random walks as being dynamic.
In this paper, we will adopt the dynamic approach by focusing on the hitting times of random walks. Given a hypergraph $H$, we select a node $t$ as target, and perform random walks starting from any other node $s$. The hitting time for this process is defined as the number of steps a random walker needs to traverse before it hits the target $t$ for the first time. By definition, the hitting time on a hypergraph is a random variable that depends on the hypergraph structure, the starting node $s$, the target node $t$, and can thus be denoted as $N^{(s)}_{t}$. The expectation of $N^{(s)}_{t}$ provides a natural measurement of the distance from $s$ to $t$. Note that the distance here is asymmetric, meaning that the distance from $s$ to $t$ is not guaranteed to be identical to that from $t$ to $s$. Since we want to use the expected hitting time to measure the closeness between two persons in the real world, and we know that the real-world human-human relationships are generally asymmetric, this lack of symmetry is thus considered desirable.
@@
The dynamics of random walks on hypergraphs are fully captured by the transition matrix, and different definitions of transition matrices produce different expected hitting times. We calculate hypergraph node distances to find the nearest neighbors of a node. To achieve this, we select a node as the target, and calculate the expected hitting times starting from all other nodes. We then rank all those nodes according to their distances with respect to the target, and select the top-$N$ nodes as the target’s nearest neighbors. If we represent real-world human-human interactions using a hypergraph, we expect to find a person’s close friends from among his/her nearest neighbors in the hypergraph. In this paper, we will evaluate the quality of our method’s results by creating a hypergraph representation of real-world human relationships and checking whether a person’s close friends show up in his/her top neighbor list.
III.1 Simple random walks on hypergraphs
@@
As already stated, the computation of hypergraph hitting times requires specification of the random walk’s transition matrix. To capture a random walker’s preference towards a more cohesive hyperedge, we generalize the method in Ref. Carletti et al. (2020) from unweighted to weighted hypergraphs, and define transition probability from node $i$ to node $j$ as
$T^{S}_{ij,i\neq j}=\frac{\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}{\sum_{k\neq i}\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{k\alpha}\}}$ (2)
@@
III.2 Insufficiencies of simple random walks
Although simple random walk as described in Section III.1 is a good foundational algorithm for exploring hypergraph properties and can be used to generate random paths for DeepWalk, we find that the expected hitting times derived from this algorithm are not a good measurement of node distances when the hypergraph is heavily weighted and scale-freeBarabási and Albert (1999). Using a hypergraph representation of characters in Harry Potter novels, the details of which data set will be given in Section IV.2, we illustrate the limitations of simple random walks.
In this hypergraph data set, Harry Potter is the central node with the highest node degree, and we want to find his close friends in this hypergraph. From our knowledge of the novels, we know that Ron Weasley and Hermione Granger are his close friends. Thus, if the expected hitting times of simple random walks is a good measurement of node distances, then Ron Weasley and Hermione Granger should have the smallest distances with respect to Harry Potter. Yet, after computing the expected hitting times of simple random walks with Harry Potter as the target node, we find that Hokey who is a house-elf has the smallest distance. The reason is that Hokey appears in the novels only once, and this single appearance coincides with that of Harry Potter. Consequently, in the hypergraph, Hokey has only a few connections, and all of these connections are directed towards Harry Potter. Therefore, a random walker starting from Hokey can rapidly hit Harry Potter. On the other hand, Hermione Granger, who is the heroine of the novels, has many connections with other characters than Harry Potter, and thus the expected hitting time from her to Harry Potter would be larger.
This observation compels us to conclude that simple random walks are insufficient for measuring node distances in highly complex hypergraphs, particularly concerning the protagonists. To accurately measure node distances, we need to design a mechanism to facilitate the transition from one protagonist to another while frustrating the transitions from minions to protagonists. This led to our proposal of the concept of frustrated random walks, which we will elaborate in the next section.
III.3 Frustrated random walks on hypergraphs
In the previous section, we demonstrated the insufficiencies of simple random walks using a hypergraph constructed from the Harry Potter novels. A key characteristic of this hypergraph is that the protagonists, though much fewer in number compared to the minions, have significantly higher node degrees. Hypergraphs with such a feature are mostly scale-free, meaning that their node degrees follow the power law distribution. Scale-free hypergraphs emerge naturally when generated following the principle of preferential attachment, which means when we introduce a new node to an existing graph or hypergraph, we attach it to a node with probability proportional to the node degree. Graphs or hypergraphs that are generated following this principle are guaranteed to be scale-freeBarabási and Albert (1999).
Preferential attachment is ubiquitous in real world, with examples including the accumulation of personal wealth, the number of times of a scientific paper being cited, the number of followers an account attracts in social media, etc. This is equivalent to the Matthew Effect, which postulates that the more one has, the easier it is for one to acquire even more. For instance, in the Harry Potter novels, a random character has a much higher probability of interacting with the protagonists than with the peripheral characters. Similarly, in academia, a highly cited researcher has a much higher probability of being cited by a new paper than an obscure one. Preferential attachment means a central node can easily be linked with lots of peripheral nodes, who would then make rapid transitions to the central node under the scenario of simple random walks. Consequently, when we use the expected hitting times of simple random walks to measure node distances, we will find that the nearest neighbors of a central node are mostly peripheral ones. We want to avoid this by introducing the concept of frustrated random walks.